A laser emits a narrow beam of light. The diameter of the beam is 3 mm, and the power is 10 mW. What is the intensity of the laser beam?

The answer is 1400 W/m^2.

Please show step by step how to get the answe to this problem showing formulas.
Answer #1

The diameter of the beam is 3 mm,

The power is 10 mW.

S0 its radious is r =1.5 mm=1.5x10-3 m

The beam area is A=\pi r^{2}=3.14x(1.5x10-3 )2 =7.065X10-6 m2

Intensity of beam I = Power/Area

I=\frac{10X10^{-3}}{7.065X10^{-6}}

I=1415 Watts/m2